The length of a certain rectangle is 4 feet greater than the width. If the length is decreased by 3 feet and the width is increased by 5 feet, the area will increased by 21 square feet. What were the dimensions of the rectangle?
Answer by VFBundy(438)

Width = w
Length = w + 4

(w + 5)((w + 4) - 3) = w(w + 4) + 21

(w + 5)(w + 1) = w² + 4w + 21

w² + 6w + 5 = w² + 4w + 21

6w + 5 = 4w + 21

2w = 16

w = 8

Width = w = 8
Length = w + 4 = 4 + 8 = 12